OpenStudy (anonymous):

1. Suppose that the amount of algae in a pond triples every 2 hours. If the pond initially contains 80 pounds of algae, how much algae will be in the pond after 8 hours? (1 point) 720 1,280 6,480 320 2. What is the order of the numbers from least to greatest? A = 1.5 × 103; B = 1.4 × 10–1; C = 2 × 104; D = 1.4 × 10–2 (1 point) D < B < A < C C < A < B < D D < B < C < A B < D < A < C

OpenStudy (anonymous):

3. Danielle has decided to raise rabbits but has been warned that the number of rabbits she will have will grow very quickly. Danielle started out with 15 rabbits and the function y = 15 • 1.8x models the number of rabbits she will have after x months. Which graph represents this function? (1 point) graph graph graph graph 4. Multiply. Give your answer in standard form. (2n2 + 5n + 4)(2n – 4) (1 point) 4n3 – 2n2 + 28n – 16 4n3 + 12n2 – 2n – 16 4n3 + 2n2 – 12n – 16 4n3 + 18n2 – 28n – 16 5. Multiply. (4x + 3)(4x – 3) (1 point) 16x2 – 24x – 9 16x2 + 24x + 9 16x2 – 9 16x2 + 9 6. The area of a rectangular painting is given by the trinomial x2 – 5x – 36. What are the possible dimensions of the painting? Use factoring. (1 point) (x + 9) and (x – 4) (x + 9) and (x + 4) (x – 9) and (x – 4) (x – 9) and (x + 4) 7. Graph the function and identify the domain and range. y = –0.25x2 (1 point) graph domain: (–∞,∞) range: y ≤ 0 graph domain: (–∞,∞) range: y ≥ 0 graph domain: (–∞,∞) range: y ≥ 0 graph domain: (–∞,∞) range: y ≤ 0 8. A catapult launches a boulder with an upward velocity of 92 m/s. The height of the boulder, h, in meters after t seconds is given by the function h = –5t2+ 92t + 16. What is the boulder’s maximum height? How long does it take to reach maximum height? Round to the nearest hundredth, if necessary. (1 point) Reaches a maximum height of 16.00 meters in 18.4 seconds. Reaches a maximum height of 18.57 meters in 9.2 seconds. Reaches a maximum height of 37.14 meters in 18.4 seconds. Reaches a maximum height of 439.20 meters in 9.2 seconds. 9. Solve the equation using the Zero Product Property. –2x(5x – 2) = 0 (1 point) x = 0, 2 over 5 x = 0, 2 x = 0, negative 2 over 5 x = 0, –2 10. Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x2 + 10 = –7x (1 point) x = 2, 5 x = –2, 5 x = 2, –5 x = –2, –5

OpenStudy (anonymous):

Help??? Thanks!! @kewlgeek555 @lauluisa @timaashorty @InYourHead @brendalopez @e.cociuba @TuringTest @esshotwired

OpenStudy (anonymous):

(I didn't want to get in trouble for posting my answers ^^)

OpenStudy (anonymous):

Did you get these?

OpenStudy (anonymous):

did any get there if so pls msg me